# CNN-JEPA: Self-Supervised Pretraining Convolutional Neural Networks Using Joint Embedding Predictive Architecture

### v1

@@ Abstract
Self-supervised learning (SSL) has become an important approach in pretraining large neural networks, enabling unprecedented scaling of model and dataset sizes. While recent advances like I-JEPA have shown promising results for Vision Transformers, adapting such methods to Convolutional Neural Networks (CNNs) presents unique challenges. In this paper, we introduce CNN-JEPA, a novel SSL method that successfully applies the joint embedding predictive architecture approach to CNNs. Our method incorporates a sparse CNN encoder to handle masked inputs, a fully convolutional predictor using depthwise separable convolutions, and an improved masking strategy. We demonstrate that CNN-JEPA outperforms I-JEPA with ViT architectures on ImageNet-100, achieving 73.3% linear top-1 accuracy with a standard ResNet-50 encoder. Compared to other CNN-based SSL methods, CNN-JEPA requires 17-35% less training time for the same number of epochs and approaches the linear and k-NN top-1 accuracies of BYOL, SimCLR, and VICReg. Our approach offers a simpler, more efficient alternative to existing SSL methods for CNNs, requiring minimal augmentations and no separate projector network.
@@ I Introduction
Self-supervised learning (SSL) became an important approach in pretraining large neural networks that allowed to scale model and dataset size to levels previously not possible. It is one of the driving mechanisms behind the recent rapid progress in large-language models. Self-supervised pretraining presents a novel paradigm which aims at learning general representations, rather than learning task-specific features.
In computer, vision designing good self-supervised pretext learning tasks is a challenging problem, with many methods proposed in the literature. One such approach is masked image modeling (MIM), where patches of the input image are masked, and the network is trained to predict the masked patches. Such methods were proposed for Vision Transformers (ViTs) primarily [1], but were also introduced for Convolutional Neural Networks (CNNs) [2, 3]. A more recent, successful approach is I-JEPA [4] (Image Joint Embedding Predictive Architecture) that essentially performs masked image modeling in the latent space by predicting the latent representation of the masked patches from provided context patches. This approach is particularly well-suited for Vision Transformers, which provide good scaling to large datasets and allow for simple implementation of masking.
Figure 1: Comparing CNN-JEPA (our method) to I-JEPA [4] and common SSL methods based on Linear top-1 accuracy and Training cost on ImageNet-100. The area of the markers are proportional to the number of parameters in the model.
Despite the growing popularity of Vision Transformers, Convolutional Neural Networks still provide competitive performance on many tasks, and are widely used in practice. They are specifically efficient and provide better performance on small to medium-sized datasets where the inductive biases of CNNs provide an advantage. However, adapting I-JEPA to CNNs is non-trivial, as neither masking nor the prediction from a dense feature map is straightforward. Computing feature maps for a masked input requires masked convolution and careful construction of the mask, that takes the downsampling of the network into account. Predicting masked patches is also challenging because it requires sufficiently large receptive field that is able to predict the latent representation based on provided context. Typical CNNs achieve large receptive fields by stacking many convolutional layers, which can be extensively parameter heavy and computationally expensive for feature maps with high depth.
In this work we propose a novel self-supervised learning method, CNN-JEPA (Convolutional Neural Network-based Joint Embedding Predictive Architecture), that addresses the challenges of adapting I-JEPA to CNNs. The masked image modeling approach proposed Tian et al. [3] uses masked convolutions and masks that are constructed based on the downsampling of the network. We introduce such sparse CNN encoder to I-JEPA to address the challenges of masking. We also introduce a novel, fully convolutional predictor architecture using depthwise separable convolutions that can predict the latent representation of the masked patches with low parameter count and computational cost. Additionally, we improve and simplify the masking strategy of I-JEPA by predicting masked patches as a single region based on the context region as the input for the predictor, instead of using multiple target regions as used by Assran et al. [4]. Our main contributions are:
@@
    A novel self-supervised learning method, CNN-JEPA, that adapts the successful I-JEPA method to Convolutional Neural Networks.
@@
    Sparse CNN encoder architecture that can correctly handle masked inputs and produce masked feature maps based on masked CNNs proposed by Tian et al. [3].
@@
Source code supporting the findings of this study is available at https://github.com/kaland313/CNN-JEPA.
@@ II Related Work
Many SSL methods have been proposed in the literature. We propose to categorize the ones that are most relevant to our work into two main groups: instance discriminative and reconstruction-based approaches. Instance discriminative methods, such as BYOL[5], SimCLR[6], and VICReg[7], learn representations by maximizing similarity between augmented views of the same image. These methods employ different mechanisms to avoid trivial solutions, including the use of negative samples, asymmetric architecture, momentum encoder, and regularization.
Reconstruction-based methods primarily rely on masked image modeling pretext tasks, where parts of the input image are masked, and the network is trained to predict the pixels of the masked regions. These methods were initially proposed for Vision Transformers[1], but were also adapted to Convolutional Neural Networks[2, 3]. Earlier works applied other reconstruction-based pretext tasks, such as predicting the color of the image from the grayscale version[8]; however, masked image modeling has shown to be more effective.
Instance discriminative methods learn image-level features that are invariant to the augmentations used to generate views, and aim to capture high-level semantics of the image. On the other hand, these methods rely heavily on hand crafted augmentations and are not trained to capture local details. In contrast, reconstruction-based methods learn more localized features that capture local detail information necessary to solve the reconstruction task; however, their pixel reconstruction objective function might not capture high-level semantics of the image. Assran et al. [4] proposed I-JEPA, a method that combines the benefits of both approaches by performing masked image modeling in the latent space, where the network is trained to predict the latent representation of the masked patches. I-JEPA was primarily designed for Vision Transformers, and adapting it to Convolutional Neural Networks has not been explored in the literature to the best of our knowledge.
@@ III Method
Figure 2: Overview of the CNN-JEPA method. The student and teacher encoders share a common architecture, such as a ResNet-50, with the student encoder using sparse convolutions. The learning objective is to predict the latent representations of masked patches using a predictor, which also trains the student encoder.
@@ III-A CNN-JEPA Algorithm
The overview of the proposed CNN-JEPA algorithm is shown on fig. 2. The proposed method trains an encoder by masking patches of the input image and predicting the latent representation of the masked patches using a predictor. The target latent representations are obtained by a teacher encoder whose weights are updated using the exponential moving average of the student encoder’s weights. We introduce a sparse CNN architecture for the student encoder that can correctly handle masked inputs and produce masked feature maps. For training, we provide the same image to both the student and teacher encoders, with masking applied only to the student encoder’s input. We refer to the masked image and features as the context, based on which the predictor produces the latent representation at the masked locations. Before feeding the masked latent representation to the predictor, masked locations in the student feature maps are filled with a learnable mask token, as proposed by Assran et al. [4]. These feature maps are then passed to a predictor, which is a CNN that predicts the latent representation of the masked patches.
The training objective is to minimize the L2 loss between the predicted latent representations and the target latent representations at the masked locations. Loss is not computed for non-masked patches, using the complement of the context-mask as the mask for the loss computation.
III-B Student and teacher encoder
The student and teacher networks in our proposed method share a common architecture, similarly to other self-supervised learning methods that employ asymmetric student-teacher mechanisms (such as BYOL [5], DINO [9] and I-JEPA [4]). However, a key distinction in our approach is that the student network utilizes sparse convolutions, as introduced by Liu et al. [10] and later employed by Tian et al. [3] for masked image modeling. This sparse convolutional architecture is necessary for handling masked inputs and producing masked feature maps. In practice, we implement sparse convolutions by masking the outputs of standard convolutional layers as detailed in section IV-A. This allows for easily updating the weights of the teacher network based on the student weights.
The teacher network employs standard convolutions. Following Assran et al. [4], the student network is trained using gradient-based optimization techniques, while the weights of the teacher network are updated through an exponential moving average of the student weights. This update mechanism ensures that the teacher network provides stable target representations for the student. We also refer to the student and teacher encoders as context and target encoders, respectively.
@@ III-C Predictor
The predictor in our proposed method operates on feature maps and predicts the latent representation of the masked patches.
Feature maps typically have high depth (e.g., 768 for ConvNeXts [11] and ViT-Base [12], 2048 for ResNet-50 [13]), but low spatial resolution (e.g., 7x7 when using a ResNet-50 on 224x224 images). As the output of the predictor must have the same dimensionality and shape as the feature maps, its outputs also have such depth and spatial size properties. The number of parameters scales quadratically with the depth for convolutional layers that have the same input and output shape. Hence, for high depth, they require a large number of parameters and are computationally expensive. To address this, we propose to use depthwise separable convolutions, which are more parameter efficient and computationally cheaper than standard convolutions. With standard convolutions, the number of parameters in the predictor can easily exceed the number of parameters in the encoder, making the training inefficient, as more resources are spent on the predictor than the encoder. Our use of depthwise separable convolutions helps mitigate this issue, ensuring a more balanced distribution of computational resources between the predictor and the encoder.
@@ III-D Mask token
The masked locations in the student feature maps are filled with a learnable mask token, before passing them to the predictor, similarly to how masked image modeling [1] and I-JEPA [4] uses a mask token. While such use of a mask token is common with vision transformers, SparK [3] also utilizes it for masked auto-encoding with CNNs. In our case, the mask token is a parameter vector with the same dimensionality as the depth of the feature maps and is shared across all masked locations and samples of a minibatch.
@@ III-E Masking
Figure 3: Example masked images, generated using multi-block and random masking strategies. The multi-block approachs masks multiple connected regions in rectangular blocks. Random masking samples patches uniformly from the image up to a given masking ratio. We employ both masking strategies, with a 0.75 probability of using multi-block masks for a given minibatch of images.
In masked image modeling, the most common masking strategy is random masking, where patches are uniformly sampled up to a given masking ratio. I-JEPA [4] introduces the multi-block masking scheme, where blocks of multiple patches are sampled, patches within blocks are masked, while patches outside the blocks form the context from which the predictor predicts the latent representations. I-JEPA presents an empirical study on random, single and multi-block masking and shows that the latter is the most effective by a large margin. We adapt the multi-block masking strategy to CNNs, however we handle the prediction of masked patches differently. While I-JEPA predicts latent features for each block separately, we consider the union of the blocks as the masked region and predict the latent representation of the whole region. For CNNs, the latter approach is simpler to implement and equivalent in terms of learning. The context for the prediction are patches outside any masked patch in both our work and I-JEPA.
When using masked image modeling with CNNs, the mask size is determined by the downsampling of the network. Typically, ResNets and ConvNeXts halve the resolution of feature maps in 5 stages (via pooling) resulting in 32 times downsampling. In order to keep the masking consistent with feature map resolution, we use masks with patch size of $32\times 3232 × 32$ pixels, which corresponds to $1\times 11 × 1$ spatial dimensions (”pixels”) for final feature maps. For images with $224\times 224224 × 224$ resolution, feature maps have $7\times 77 × 7$ spatial dimensions, with matching $7\times 77 × 7$ masks of $32\times 3232 × 32$ pixel patches. For ViTs, the patch size is determined by the patch size of the transformer, which is typically $16\times 1616 × 16$ or $14\times 1414 × 14$ pixels.
@@ III-F Comparison to other methods
Compared to common instance discriminative SSL methods one key advantage of our proposed method is its simplicity. It requires minimal handcrafted augmentations, relying only on random resized crop and masking.
Additionally, our approach doesn’t require a separate projector network as opposed to most SSL methods, further simplifying the architecture and reducing computational overhead. Many SSL approaches learn features invariant to diverse augmentations, necessitating a projector network after the encoder. The use of such projector allow the encoder to learn more general features, while the projector learns mapping specific to the loss function and augmentations of the SSL method. Similarly to I-JEPA [4], our approach learns general features without the need for such projector network.
@@ IV Experiments
IV-A Encoder
We conduct experiments using a ResNet-50[13] encoder, for its common use in SSL literature. Sparse convolutions are implemented by zeroing out the outputs of each convolutional layer using a mask appropriately upscaled to the layer’s activation dimensions. While this is computationally inefficient, it is simple to implement and can use standard convolutional layers, which are optimized for many GPU architectures. Using sparse convolutional layers would require a specialized GPU library (e.g., Minkowski Engine [14]), and it would make the implementation of our method more complex and harder to reproduce. Additionally, using standard convolutions simplify the exponential moving average update of the teacher encoder’s weights.
IV-B Predictor
@@
IV-C Datasets
We conduct experiments on ImageNet-100 and ImageNet-1k [16] datasets. With encoder architectures that downscale by a factor of 32, such as a ResNet-50, the patch size must be at least 32x32 pixels, therefore images must be sufficiently larger than 32x32 pixels to provide meaningful target and context regions for the masked embedding prediction task. Hence, we use 224x224 pixel images from ImageNet-1k and ImageNet-100. Additionally, for these reasons, we do not conduct experiments with lower resolution images and datasets such as CIFAR-10 or STL-10.
IV-D Classification Probes
To evaluate the quality of the learned representations, we use linear and k-nearest neighbor (k-NN) classification probes on top of the encoder’s frozen features. We train the fully connected layer of the linear probe for 90 epochs and report the best top-1 and top-5 accuracies. For k-NN classification, we use the same features and also report top-1 and top-5 accuracies. For both probes, we use minimal augmentations: deterministic resizing and cropping, and normalization. We adapt the hyperparameters of both probes from Lightly [17].
TABLE I: Linear and k-NN top-1 accuracies for different algorithms and backbones on ImageNet-1k and ImageNet-100 datasets. All experiments were limited to 100 epochs for ImageNet-1k and 200 epochs for ImageNet-100.
@@ IV-D Classification Probes
Linear top-1 k-NN top-1 Linear top-1 k-NN top-1
Backbone [# Paramters] Algorithm
ResNet-50 [23.5 M] SimCLR [6] 76.20 67.26 59.81 45.31
@@
ViT-Small [21.6 M] I-JEPA[4] 42.30 34.84 - -
ViT-Base [85.7 M] 46.36 31.36 - -
IV-E Training Details
We use the AdamW optimizer with a constant weight decay of 0.01. A cosine learning rate scheduler with warm-up is employed for 10 epochs, reaching a peak learning rate of 0.01. The per-device batch size is 128 with 4 GPUs, resulting in a total batch size of 512. We update the weights of the teacher encoder using the exponential moving average approach common in other SSL methods [5, 4], with the momentum parameter increasing from 0.996 to 1.0 over the full training period.
Training is conducted for 200 epochs on ImageNet-100 and 100 epochs on ImageNet-1k. All pretraining runs are performed on 4 NVIDIA A100 40GB GPUs on a cluster, with each node having 4 GPUs, 256 GB RAM, and 64 CPU cores. Pretraining on ImageNet-100 for 200 epochs on this hardware takes 13 hours, while pretraining on ImageNet-1k for 100 epochs takes 70 hours.
We utilize PyTorch, PyTorch Lightning [18], and the Lightly [17] SSL library for implementing baselines, probes, and our method. We obtain the I-JEPA ViT results from the official implementation [4].
The cumulative runtime for all experiments in this research is 10 000 hours with an estimated total emission of 1000 kgCO2eq [^1] Estimations were conducted using the MachineLearning Impact calculator presented in [19].

### v2

@@ Abstract
Self-supervised learning (SSL) has become an important approach in pretraining large neural networks, enabling unprecedented scaling of model and dataset sizes. While recent advances like I-JEPA have shown promising results for Vision Transformers, adapting such methods to Convolutional Neural Networks (CNNs) presents unique challenges. In this paper, we introduce CNN-JEPA, a novel SSL method that successfully applies the joint embedding predictive architecture approach to CNNs. Our method incorporates a sparse CNN encoder to handle masked inputs, a fully convolutional predictor using depthwise separable convolutions, and an improved masking strategy. We demonstrate that CNN-JEPA outperforms I-JEPA with ViT architectures on ImageNet-100, achieving a 73.3% linear top-1 accuracy using a standard ResNet-50 encoder. Compared to other CNN-based SSL methods, CNN-JEPA requires 17-35% less training time for the same number of epochs and approaches the linear and k-NN top-1 accuracies of BYOL, SimCLR, and VICReg. Our approach offers a simpler, more efficient alternative to existing SSL methods for CNNs, requiring minimal augmentations and no separate projector network.
@@ I Introduction
Self-supervised learning (SSL) became an important approach in pretraining large neural networks, allowing model and dataset sizes to scale to levels that were previously unattainable. In computer vision, designing good self-supervised pretext learning tasks is a challenging problem, with many methods proposed in the literature. One such approach is I-JEPA [1] (Image Joint Embedding Predictive Architecture), which learns by predicting the latent representation of the masked patches from latent representations of provided context patches. This approach is particularly well-suited for Vision Transformers. However, adapting I-JEPA to Convolutional Neural Networks is non-trivial, as neither masking nor predicting from a dense feature map is straightforward. Computing feature maps for a masked input requires masked convolution and careful construction of the mask, that takes the network’s downsampling into account. Predicting masked patches is also challenging because it requires a sufficiently large receptive field capable of predicting the latent representation based on the provided context. Typical CNNs achieve large receptive fields by stacking many convolutional layers, which can be extensively parameter-heavy and computationally expensive for feature maps with high depth.
Figure 1: Comparing CNN-JEPA (our method) to I-JEPA [1] and common SSL methods based on linear top-1 accuracy, and training cost on ImageNet-100. The area of the markers is proportional to the number of parameters in the model.
In this work, we propose a novel self-supervised learning method, which we call CNN-JEPA (Convolutional Neural Network-based Joint Embedding Predictive Architecture), that addresses the challenges of adapting I-JEPA to CNNs. The masked image modeling approach proposed by Tian et al. [2] uses masked convolutions and masks that are constructed based on the downsampling of the network. We introduce such sparse CNN encoder to I-JEPA to address the challenges of masking. We also introduce a novel, fully convolutional predictor architecture using depthwise separable convolutions that can predict the latent representation of the masked patches with low parameter count and computational cost. Additionally, we improve and simplify the masking strategy of I-JEPA [1] by predicting masked patches as a single region.
Our main contributions are:
@@
    A novel self-supervised learning method, CNN-JEPA that adapts the successful I-JEPA method to Convolutional Neural Networks.
@@
    Sparse CNN encoder architecture that can correctly handle masked inputs and produce masked feature maps, building on the masked CNNs proposed by Tian et al. [2].
@@
The source code supporting the findings of this study is available at https://github.com/kaland313/CNN-JEPA.
@@ II Related Work
We propose to categorize SSL methods that are most relevant to our work into two main groups: instance discriminative and reconstruction-based approaches. Instance discriminative methods, such as BYOL [3], SimCLR [4], and VICReg [5], learn representations by maximizing similarity between augmented views of the same image.
Reconstruction-based methods primarily rely on masked image modeling pretext tasks, where parts of the input image are masked, and the network is trained to predict the pixels of the masked regions. These methods were initially proposed for Vision Transformers [6], but were also adapted to Convolutional Neural Networks [2].
Instance discriminative methods learn image-level features that are invariant to the augmentations used to generate views, and aim to capture high-level semantics of the image.In contrast, reconstruction-based methods learn more localized features that capture detailed information necessary to solve the reconstruction task; however, their pixel reconstruction objective function might not capture the high-level semantics of the image. Assran et al. [1] proposed I-JEPA, a method that combines the benefits of both approaches by performing masked image modeling in the latent space, where the network is trained to predict the latent representation of the masked patches. I-JEPA was primarily designed for Vision Transformers, and adapting it to Convolutional Neural Networks has not been explored in the literature to the best of our knowledge.
@@ III Method
Figure 2: Overview of the CNN-JEPA method. The context and target encoders share a common architecture, with the context encoder using sparse convolutions. The learning objective is to predict the latent representations of masked patches using a predictor, which also trains the context encoder.
@@ III-A CNN-JEPA Algorithm
The overview of the proposed CNN-JEPA algorithm is shown on Fig. 2. The proposed method trains an encoder by masking patches of the input image and predicting the latent representation of the masked patches using a predictor. The target latent representations are obtained by a target encoder whose weights are updated using the exponential moving average of the context encoder’s weights. We introduce a sparse CNN architecture for the context encoder that can correctly handle masked inputs and produce masked feature maps. For training, we provide the same image to both the context and target encoders, with masking applied only to the context encoder’s input. We refer to the masked image and features as the context, based on which the predictor produces the latent representations at the masked locations. Before feeding the masked latent representation to the predictor, masked locations in the context feature maps are filled with a learnable mask token.These feature maps are then passed to a predictor, which is a CNN designed to predict the latent representation of the masked patches.
The training objective is to minimize the L2 loss between the predicted and target latent representations at the masked locations. Loss is not computed for non-masked patches.
III-B Context and target encoder
The context and target networks in our proposed method share a common architecture. The context network utilizes sparse convolutions, as introduced by Liu et al. [7] and later employed by Tian et al. [2] for masked image modeling. This sparse convolutional architecture is necessary for handling masked inputs and producing masked feature maps. The target network uses standard convolutions.
The context network is trained using gradient-based optimization, while the target network’s weights are updated using an exponential moving average of the context weights. This update mechanism ensures that the target network provides stable target representations for the loss function. In practice, we implement sparse convolutions by selectively masking the outputs of standard convolutional layers. This allows for easily updating the weights of the target network based on the context weights.
@@ III-C Predictor
The predictor in our proposed method operates on feature maps and predicts the latent representation of the masked patches. Feature maps typically have high depth (e.g., 2048 for ResNet-50 [8] and 768 ViT-Base [9]), but low spatial resolution. As the output of the predictor must have the same dimensionality and shape as the feature maps, its outputs also have such depth and spatial size properties. The number of parameters scales quadratically with the depth for convolutional layers that have the same input and output shape. Hence, for high depth, they require a large number of parameters and are computationally expensive. To address this, we propose to use depthwise separable convolutions, which are more parameter efficient and computationally cheaper than standard convolutions. With standard convolutions, the number of parameters in the predictor can easily exceed the number of parameters in the encoder. Our use of depthwise separable convolutions helps mitigate this issue, ensuring a more balanced distribution of computational resources between the predictor and the encoder.
@@ III-D Mask token
The masked locations in the context feature maps are filled with a learnable mask token, before passing them to the predictor, similarly to how masked image modeling [6, 2] and I-JEPA [1] use a mask token. In our case, the mask token is a parameter vector with the same dimensionality as the depth of the feature maps and is shared across all masked locations and samples of a minibatch.
@@ III-E Masking
In masked image modeling, the most common masking strategy is random masking, where patches are uniformly sampled up to a given masking ratio. I-JEPA [1] introduces the multi-block masking scheme, where blocks of multiple patches are sampled, patches within blocks are masked, while patches outside the blocks form the context from which the predictor predicts the latent representations. I-JEPA presents an empirical study on random, single, and multi-block masking and shows that the latter is the most effective by a large margin. We adapt the multi-block masking strategy to CNNs. We consider the union of the blocks as the masked region and predict the latent representations in it. The context for the prediction is patches outside any masked patch.
When using masked image modeling with CNNs, the mask size is determined by the downsampling of the network. Typically, ResNets halve the resolution of feature maps in 5 stages (via pooling), resulting in 32 times downsampling. In order to keep the masking consistent with feature map resolution, we use masks with a patch size of $32\times 3232 × 32$ pixels, which corresponds to $1\times 11 × 1$ spatial dimensions (’pixels’) for final feature maps.
@@ III-F Comparison to other methods
Compared to common instance discriminative SSL methods, one key advantage of our proposed method is its simplicity. It requires minimal handcrafted augmentations, relying only on random resized crop and masking.
Additionally, our approach doesn’t require a separate projector network as opposed to most SSL methods, further simplifying the architecture and reducing computational overhead. Many SSL approaches learn features invariant to diverse augmentations, necessitating a projector network after the encoder. Our approach learns general features without the need for such a projector network.
@@ IV Experiments
IV-A Encoder and predictor
We conduct experiments using a ResNet-50 [8] encoder for its common use in SSL literature. Sparse convolutions are implemented by zeroing out the outputs of each convolutional layer using a mask appropriately upscaled to the layer’s activation dimensions. While this is computationally inefficient, it is simple to implement and can use standard convolutional layers, which are optimized for many GPU architectures. Using sparse convolutional layers would require specialized GPU code, making the implementation of our method more complex and harder to reproduce.
@@ IV-A Encoder and predictor
IV-B Classification Probes
To evaluate the quality of the learned representations, we use linear and k-nearest neighbor (k-NN) classification probes on top of the encoder’s frozen features. We train the fully connected layer of the linear probe for 90 epochs and report the best top-1 and top-5 accuracies. For k-NN classification, we use the same features and also report top-1 and top-5 accuracies. For both probes, we use minimal augmentations: deterministic resizing and cropping, and normalization.
TABLE I: Linear and k-NN accuracies on ImageNet-1k and ImageNet-100
@@ IV-B Classification Probes
Linear k-NN Linear k-NN
Backbone Algorithm top-1 top-1 top-1 top-1
ResNet-50 SimCLR [4] 76.20 67.26 59.81 45.31
@@
ViT-Small I-JEPA[1] 42.30 34.84 -a
ViT-Base 46.36 31.36
  • a
    Assran et al. [1] only published ImageNet-1k results for larger ViT models, which would not provide a fair comparison to ResNet-50. Additionally, we were unable to obtain results for the ViT-Small and ViT-Base models using their official code on our hardware.
IV-C Training Details
We conduct experiments on ImageNet-100 and ImageNet-1k [10] datasets. We use a patch size of $32\times 3232 × 32$ pixels and $224\times 224224 × 224$ pixels resolution. We use the AdamW optimizer with a constant weight decay of 0.01. A cosine learning rate scheduler with warm-up is employed for 10 epochs, reaching a peak learning rate of 0.01. The per-device batch size is 128 with 4 GPUs. We update the target encoder’s weights using the exponential moving average approach common in SSL methods [3, 1], with the momentum increasing from 0.996 to 1.0 during training. All pretraining runs use 4 NVIDIA A100 40GB GPUs, 256 GB RAM and 64 CPU cores. Pretraining is conducted for 200 epochs on ImageNet-100 and 100 epochs on ImageNet-1k, taking 13 hours and 70 hours, respectively.